OpenStudy (anonymous):

. Can the sum of two polygons in x, both of degree 4, be of degree 3? If so give an example. Can somebody help me with this please ????????

OpenStudy (anonymous):

i assume by polygons you meant polynomials; i will give a simple case x^4 + x^4 = 2x^4 two polynomials of degree 4, = polynomial of degree 4 however, what about this? (x^4 + x^3) + (x^4) = 2x^4 + x^3 Two polynomials, highest degree 4, equals a polynomial of degree 4 If we keep adding polynomials together, we will never get a lower degree than 4. However, if we subtract the polynomials such as this: (x^4 + x^3) - (x^4) = x^3 We will get a lower degree. But that is not allowed per your question. Thus, as long as we add two polynomials of degree 4, we will always get two polynomials of at least degree 4.

OpenStudy (anonymous):

thankss and yes i meant polynomials thanks so much